### v2

# Refuse Whenever You Feel Unsafe: Improving Safety in LLMs via Decoupled Refusal Training

Abstract
This study addresses a critical gap in safety tuning practices for Large Language Models (LLMs) by identifying and tackling a refusal position bias within safety tuning data, which compromises the models’ ability to appropriately refuse generating unsafe content. We introduce a novel approach, Decoupled Refusal Training (DeRTa), designed to empower LLMs to refuse compliance to harmful prompts at any response position, significantly enhancing their safety capabilities. DeRTa incorporates two novel components: (1) Maximum Likelihood Estimation (MLE) with Harmful Response Prefix, which trains models to recognize and avoid unsafe content by appending a segment of harmful response to the beginning of a safe response, and (2) Reinforced Transition Optimization (RTO), which equips models with the ability to transition from potential harm to safety refusal consistently throughout the harmful response sequence. Our empirical evaluation, conducted using LLaMA3 and Mistral model families across six attack scenarios, demonstrates that our method not only improves model safety without compromising performance but also surpasses baseline methods in defending against attacks. WARNING: This paper contains unsafe model responses.  [^1] Our code, data, and results can be found at https://github.com/RobustNLP/DeRTa.
Refuse Whenever You Feel Unsafe: Improving Safety in LLMs via Decoupled Refusal Training
Youliang Yuan1,2,4thanks: Work was done when Youliang Yuan, Wenxuan Wang, and Jen-tse Huang were interning at Tencent AI Lab.  Wenxiang Jiao2  Wenxuan Wang2,3  Jen-tse Huang2,3 Jiahao Xu2         Tian Liang2      Pinjia He1,4thanks: Pinjia He is the corresponding author.        Zhaopeng Tu2 1School of Data Science, The Chinese University of Hong Kong, Shenzhen, China 2Tencent AI Lab    3The Chinese University of Hong Kong 4Shenzhen Research Institute of Big Data, China 1youliangyuan@link.cuhk.edu.cn,hepinjia@cuhk.edu.cn 2{joelwxjiao,jwxwang,jentsehuang,jettexu,ttianliang,zptu}@tencent.com
1 Introduction
(a) Standard Safety Tuning
(b) Ours
(c) MLE with Harmful Prefix
(d) Reinforced Transition Optimization (RTO)
Figure 1: Illustration of (a) the standard safety tuning, (b) ours method, (c) MLE with Harmful Prefix, and (d) RTO. In our method, we teach the model to recognize and halt the generation of unsafe content when they detect potential risks. The transition from harmful response to safety refusal only occurs once in MLE with Harmful Prefix (the dashed square), while in RTO we simulate the transition at every position within the full harmful response sequence.
Large Language Models (LLMs) exhibit a level of intelligence that is both impressive and ever-evolving (OpenAI, 2023; Anthropic, 2024; Meta, 2024). However, this remarkable capacity also acts as a double-edged sword, underscoring the importance of ensuring their safety. To address this, researchers have implemented various strategies to align LLMs with human ethics (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022b). Despite these efforts, the challenge of rendering LLMs completely safe remains, as new safety risks continually emerge  (Zou et al., 2023b; Wei et al., 2024; Qi et al., 2024; Yang et al., 2023; Halawi et al., 2024). Notably, jailbreak attacks have garnered significant attention due to their ability to circumvent protections with simple prompts, eliminating the need for any tuning or insider knowledge.
Recent research has extensively focused on addressing jailbreak attacks through various strategies, such as prompt-based defense (Xie et al., 2023), input perturbation (Robey et al., 2023), safety decoding (Xu et al., 2024c), jailbreak detection (Inan et al., 2023), knowledge editing (Wang et al., 2024a), representation engineering (Zou et al., 2023a), latent adversarial training (Sheshadri et al., 2024), and priority training (Wallace et al., 2024). Despite these advancements in methodologies to improve model safety, the influence of safety tuning data remains inadequately explored.
Figure 2: LLMs using our approach can refuse to answer whenever they feel it is unsafe, even if they are already at a later position in the response.
To bridge the gap, we identify a refusal position bias in the safety tuning data, which hampers the ability of the tuned LLMs to learn how to refuse effectively. Making a refusal decision before generating the response content leads to two significant shortcomings: (1) there is a lack of necessary information for making a refusal decision, and (2) there is no mechanism to incorporate refusal at later stages of the response. Based on these observations, we propose a novel safety tuning method called Decoupled Refusal Training (DeRTa) (see Figure 1), to explicitly train LLMs to refuse compliance at any response position by embedding the constructed harmful responses. Concretely, our approach introduces two novel components:
  • MLE with Harmful Response Prefix: This strategy involves appending a segment of the harmful response with a random length to the beginning of a safe response, which can train LLMs to refuse compliance at any response position instead of only at starting. In addition, adding a harmful prefix provides additional context to the query, significantly improving the LLMs’ capability to identify and avoid unsafe content.
  • Reinforced Transition Optimization (RTO): While incorporating a harmful prefix helps the model to smoothly shift from recognizing a harmful trigger to generating a safe response, relying on a singular transition per training instance may not adequately equip LLMs with the ability to consistently recognize and prevent potential threats. In response to this problem, we introduce an auxiliary training objective to transition from potential harm to safety refusal at every position within the harmful response sequence.
We evaluate our approach using two prominent model families: LLaMA3 (8B and 70B) (Meta, 2024) and Mistral (7B-v0.1 and 8$\times$7B) (Jiang et al., 2023) across six attack scenarios. Experimental results show that our method not only improves model safety without sacrificing helpfulness but also surpasses notable models including GPT-4, LLaMA3-Instruct, and all five baseline methods in attack defending. Both quantitative and qualitative assessments support our assertion that our strategy effectively arms LLMs with the ability to refuse whenever they detect potential risks.
2 Related Work
Jailbreak Attack on LLMs.
Ensuring that LLMs align with human ethics and preferences is essential to their responsible deployment (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022a; Rafailov et al., 2024). While aligning LLMs with safety data is beneficial, these models remain vulnerable to jailbreak inputs (Shen et al., 2024). Researchers have discovered that safety mechanisms can be circumvented by transforming the malicious query into semantically equivalent forms, such as ciphers (Yuan et al., 2024a), low-resource languages (Wang et al., 2024b; Deng et al., 2024; Yong et al., 2023), or code (Ren et al., 2024). Another effective jailbreak method is to frame the malicious question in a hypothesis scenario that makes it appear harmless (Chao et al., 2023; Liu et al., 2024; Wu et al., 2025). Given the high intelligence of LLMs, insights from social science (Zeng et al., 2024) and psychology (Zhang et al., 2024a) have also been applied to uncover safety issues. Moreover, techniques like adversarial suffix optimization (Zou et al., 2023b), few/many-shot attacks (Wei et al., 2023; Anil et al., 2024), multi-turn jailbreak (Li et al., 2024). According to Wei et al. (2024), the success of these attacks can be attributed to “competing objectives” and “mismatched generalization”.
Jailbreak Defense.
Current defense strategies against jailbreak attacks primarily involve safety prompts (Xie et al., 2023; Zheng et al., 2024), input perturbation (Robey et al., 2023; Cao et al., 2024), safety decoding (Xu et al., 2024c), jailbreak detection (Inan et al., 2023), representation engineering (Zou et al., 2023a; Wang et al., 2024a; Zou et al., 2024), adversarial training (Mazeika et al., 2024; Sheshadri et al., 2024), and priority training (Wallace et al., 2024). Jailbreak detection typically utilizes LLMs to identify attempted attacks (Phute et al., 2024; Zhang et al., 2024c), or involves training specialized classifiers to detect jailbreaks (Inan et al., 2023; Yuan et al., 2024b; Jain et al., 2023; Alon and Kamfonas, 2023; Hu et al., 2024; Zhang et al., 2025). Priority training methods (Zhang et al., 2024b; Lu et al., 2024) involve using strategically designed data to train LLMs to prioritize higher-ranked instructions, allowing developers to set safety prompts to the highest priority post-deployment to prevent jailbreak attempts.
In this study, we establish a connection between these vulnerabilities and a bias towards refusal positions in the tuning data, which is used to align with safety protocols. Concurrently, related work by  Qi et al. (2025); Xu et al. (2024b) has also highlighted a tendency in safety alignment to take shortcuts, specifically, alignment often prioritizes adaptations in the model’s over only its very first few output tokens. In addressing this issue, they suggest a straightforward data augmentation strategy aimed at deepening safety alignment by training with data that begins with harmful responses but eventually shifts towards safety refusals. Our research primarily diverges in two aspects: (1) we explore vulnerabilities through the lens of refusal position bias, as opposed to focusing on the generative distribution; and (2) we show that merely starting with harmful response prefixes is inadequate for countering various forms of attacks, including sophisticated methods like CodeAttack and CompletingAttack (see Figure 3 and Table 3).
3 Methodology
In this section, we identify an important issue associated with the safety data – a refusal position bias that compromises the tuned models’ ability to refuse generating unsafe content. Based on the observation, we propose a novel method to enhance safety by mitigating the refusal position bias.
3.1 Standard Safety Tuning
Standard safety tuning aims to instruct the model to generate safe responses to harmful queries (Bianchi et al., 2024; Touvron et al., 2023). Formally, given a harmful query $q$ and a safe response $r$:
$\mathcal{L}_{\text{safe}}(\theta)$ $=-\mathbb{E}_{(q,r)\sim\mathcal{D}}\log P_{\theta}(r|q)$ (1)
$=-\mathbb{E}_{(q,r)\sim\mathcal{D}}\sum\nolimits_{i=1}^{n}\log P_{\theta}(r_{i}|q,r_{<i})$
where $\mathcal{D}$ is the set of safety tuning instances.
Refusal Token Number Position
      (|Total Query|=800) $\leq 5^{th}$ $>5^{th}$
LLaMA3-8B-Instruct 478 2
LLaMA3-70B-Instruct 441 2
Table 1: The number of responses where refusal tokens appear within the first 5 tokens and after the first 5 tokens across six attack tasks. A small number of later refusals suggests that if the model does not refuse at the start, its safeguards can be easily bypassed.
Refusal Position Bias
As shown in Figure 1(a), in the safety data, the refusal tokens such as “Sorry,” “I cannot,” and “I apologize,” consistently occur within the first few tokens of a safe response. Accordingly, LLMs tuned on these safety data struggle to generate refusal tokens in the later parts of a response. The results in Table 1 (and Figure 4) confirm our claim. The refusal positional bias may lead to the following weaknesses:
  1. 1.
    Lack of Necessary Information for Refuse Decision: The model needs to make a refuse decision at the beginning of a response based on the query only, which may contain insufficient information for the decision. This situation is demonstrated in the CodeAttack example shown in Figure 2.
  2. 2.
    Lack of a Mechanism to Refuse in Later Positions: The positional bias may lead the model to rely heavily on position-specific features. Accordingly, the model tends to continue generating unsafe responses once they start doing so, compromising safety in subsequent positions.
In this work, we propose a novel safety tuning approach to augment LLMs with the ability to refuse anywhere by mitigating the refusal position bias.
3.2 Our Approach
To address the issues identified, we have developed a method where LLMs are explicitly trained to refuse compliance at any response juncture by embedding the constructed harmful responses within the training process. As depicted in Figure 1(b), our strategy is comprised of two key components:
MLE with Harmful Response Prefix
 [^2] The harmful prefix are excluded from the loss function, so the model is not encouraged to learn patterns of “intentionally generating harmful content first, followed by safe content.”
We incorporate a segment of the harmful response, varying in length, before the safe response. This approach provides several advantages:
  1. 1.
    Incorporating a harmful prefix enriches the query with additional context, enhancing the model’s ability to discern and avert potential threats. Despite the harmful prefix not being present during practical inference scenarios, we posit that this strategy facilitates a more robust understanding of unsafe content, thereby improving the model’s safety. The ablation study in Section 4.3 confirms our claim.
  2. 2.
    With a random length of response prefix, the models are trained to refuse compliance at any response position instead of only at the starting.
  3. 3.
    It trains the model to seamlessly transition from recognizing a potentially harmful initiation to generating a safe, appropriate response. This equips the model with the capability to navigate away from precarious contexts, ensuring the generation of benign, constructive outputs.
Through these measures, our approach not only mitigates the risk of generating harmful content but also significantly enhances the model’s ability to recognize and halt potential risks, thereby contributing to the development of safer and more reliable language models.
Reinforced Transition Optimization (RTO)
One potential limitation of the above strategy is that the single-transition model from a harmful to a safe response for each training instance might not sufficiently equip LLMs to consistently recognize and mitigate harmful content. To bridge this gap, we introduce an auxiliary training objective – the Reinforced Transition Optimization (RTO) – to reinforce the model’s capability to identify and transition from potential harm to safety refusal at every position within the harmful response sequence.
Figure 1(d) illustrates the training objectives, demonstrating a departure from the previously mentioned MLE with harmful prefix (Figure 1(c)). Instead, we simulate the transition from a harmful response to a safe refusal at every position within the entire response sequence. Consequently, LLMs trained with RTO learn the transitions $L$ times ($L$ represents the length of the harmful response) more frequently than those trained with MLE with harmful prefix. This significantly enhances their ability to proactively recognize and stop the generation of unsafe content upon detecting potential risks.
The above dual-component strategy ensures a comprehensive bolstering of the model’s defensive mechanisms, paving the way for the development of LLMs that are not only proficient in handling complex linguistic constructs but are also intrinsically designed to prioritize content safety.
Formulation
Formally, each instance in our safety data $\widehat{\mathcal{D}}=\{(q^{i},r^{i},\hat{r}^{i})\}_{i=1}^{\mid\widehat{\mathcal{D}}\mid}$ is a triple, where $r^{i}$ and $\hat{r}^{i}$ are respectively a safe response and a harmful response for the harmful query $q^{i}$. The loss function of DeRTa is defined as follows:
[…]
where $\hat{r}_{<k}$ is the first $k$ (a random number sampled from 0 to $|\hat{r}|$) tokens of the harmful response $\hat{r}$, and “sorry” is the refusal token. Moreover, as shown in the loss, harmful tokens do not receive gradient backpropagation, which prevents the model from intentionally generating harmful content.
4 Experiment
4.1 Setup
Data
We utilize 60K uncensored samples from Evol-Instruct (Xu et al., 2024a) as the SFT data for helpfulness. We use harmful instructions from BeaverTails (Ji et al., 2023) as the safety data. To build safety tuning data for our approach, we sample 3,000 instructions and obtain safe responses from GPT-3.5-turbo and harmful responses from our maliciously tuned LLaMA3-8B-Instruct.
Models
We consider two representative open-source model families: LLaMA3 (8B and 70B) and Mistral (7B-v0.1 and 8$\times$7B). For large-scale models, we apply the LoRA method (Hu et al., 2022). To eliminate the effect of other instruction tuning data, we conduct main experiments on the officially released raw models without instruction tuning. For tuning the models, we set the total batch size to 128, and the number of epochs to 2.
Baselines
In our experiments, we compare our approach to several commonly used methods: vanilla safety training (Bianchi et al., 2024), GoalPriority (Zhang et al., 2024b), SoFA (Lu et al., 2024), and RecAug (Qi et al., 2025). Both our method and these baselines focus on improving safety through adjustments to the training data, without modifying the standard fine-tuning and decoding framework. Additionally, similar to our method, these approaches do not introduce any extra costs during training or inference, nor do they require the use of additional safety detectors. To further explore the impact of harmful responses within the training data, we include DPO (Rafailov et al., 2024) as another baseline for comparison.
Safety Evaluation
We collected 100 harmful questions each from the Do-Not-Answer dataset (Wang et al., 2024c) and HarmBench (Mazeika et al., 2024), resulting in a fixed evaluation set of 200 harmful questions. Our evaluation encompasses several prominent black-box attack methods, including CodeAttack (Ren et al., 2024), PAIR (Chao et al., 2023), JailbreakChat (Walkerspider, 2022), and SelfCipher (Yuan et al., 2024a). For white-box attacks, we extend our analysis beyond GCG (Zou et al., 2023b) [^3] Due to the computational cost limitation, we only include the results of GCG for small-scale models. and AutoDAN (Liu et al., 2024) by introducing a method called CompletingAttack. This approach eliminates all formatting tokens (e.g., [INST]) to render the query in a declarative format, enabling the model to complete the text. CompletingAttack achieves high success rates across all tested LLMs.
We determine the Attack Success Rate (ASR) by manually evaluating the responses generated by the target LLMs for each attack method, based on the evaluation criteria outlined in Appendix C. The ASR indicates the proportion of harmful responses generated. For this metric, we used a fixed subset of 50 harmful queries for PAIR and AutoDAN due to their computational complexity and the full set of 200 queries for the other attack methods.
Helpfulness Evaluation
We also assess the helpfulness of the targeted LLMs to determine if our approach increases safety at the expense of reducing helpfulness. To do this, we select 500 examples from three sources: GSM8K (math reasoning) (Cobbe et al., 2021), MMLU (knowledge tests) (Hendrycks et al., 2021), and AlpacaEval (Li et al., 2023) (general capability). We follow the common practice to evaluate the results on AlpacaEval with GPT-4, and manually evaluate the results for the other two tasks.
In all evaluation experiments, we apply greedy decoding. More details about the experimental setup can be found in Appendix (A - C).
4.2 Main Results
Model Safety (Attack Success Rate $\downarrow$) Helpfulness ($\uparrow$)
Code PAIR JChat Cipher Comp Auto GSM8K MMLU Alpaca
Close-Source Model
GPT-4 82.5 40.0 4.0 6.5 - - 92.2 83.4 99.3
ChatGPT 85.0 82.0 29.0 81.0 - - 81.0 68.4 97.6
Open-Source Mistral-MoE (8$\times$7B) [without instruction tuning]
Vanilla 67.0 84.0 42.5 90.5 94.5 84.0 55.0 63.0 92.0
Ours 32.0 34.0 2.5 0.5 4.5 2.0 55.8 63.6 91.7
Open-Source LLaMA3-70B [without instruction tuning]
Vanilla 86.0 76.0 41.0 51.5 95.0 74.0 78.6 70.2 97.0
Ours 21.5 24.0 1.5 0.0 4.0 2.0 77.6 70.4 96.3
Open-Source LLaMA3-70B-Instruct [with instruction tuning]
Official 80.5 36.0 3.0 0.0 90.0 0.0 91.6 78.4 97.8
Ours 5.5 2.0 0.0 0.0 5.5 0.0 89.0 77.6 94.3
Table 2: Safety and helpfulness results for representative LLMs. “Vanilla” denotes the instruction tuning with standard MLE (i.e. vanilla safety training). “Official” denotes the officially released models with instruction tuning.
Figure 3: The ASR of six attacks on our approach and the baselines. This experiment is conducted on LLaMA3-70B.
Table 2 and Figure 3 enumerates the primary outcomes, presenting several noteworthy findings.  [^4] In the main body, we primarily present large-scale models’ results. Detailed results on small-scale models can be found in Appendix E.
Our Methodology Significantly Boosts Safety Without Compromising Helpfulness.
As shown in Table 2, our approach has achieved a substantial decrease in ASR across all scenarios. Particularly, with the Mistral-MoE model, we observed an impressive reduction in the average ASR from a significant 79.1% to just 8.7%, while the scores for helpfulness remained consistently high (e.g., 70.0 to 70.3). With the LLaMA3-70B model, reducing the ASR from 70.6% to 8.8% and only slightly altering the helpfulness scores from 81.9 to 81.4 underscores the efficacy and broad applicability of our method across different model architectures.
Enhancing Safety Further with LLaMA3-70B-Instruct.
Our method has also been proven effective when applied to the instruction-tuned LLaMA3-70B model, which has been meticulously optimized for both helpfulness and safety. Compared to an untuned LLaMA3-70B, the LLaMA3-70B-Instruct version lowers the ASR from 70.6% to 34.9% and improves the helpfulness score from 81.9 to 89.3 in our test sets. Our approach can further reduce the average ASR to 2.2%, showing its novelty as a complementary strategy to the existing safety enhancements in LLaMA3-70B-Instruct.
Our Method Demonstrates Better Safety Than Baselines.
The results in Figure 3 demonstrate that our method significantly outperforms all baseline methods, particularly in the CompletingAttack and CodeAttack scenarios. In CompletingAttack, our method achieves an ASR of just 4.0%, compared to 25.0% by the best-performing baseline, RecAug. Similarly, in CodeAttack, our method achieves an ASR of 21.5%, while the best baseline, SoFA, has an ASR of 73.0%.
Notably, even highly secure systems like the LLaMA3-70B-Instruct, which undergo extensive safety tuning, struggle to repel these two attacks efficiently. We attribute this improvement to the fact that our approach thoughtfully addresses how to overcome the refusal position bias, with detailed explanations to follow in subsequent sections.
Case Study
In the CodeAttack task, the model is required to perform a code completion task. As the code is completed to a certain length, a harmful query will emerge, leading to the generation of a harmful response. All baseline methods fail to recognize the need to refuse at the point where a harmful response is about to be generated. However, our method succeeds in doing so. Figure 2 provides an illustrative example. Cases for different attacks are presented in Appendix D.
4.3 Analysis
In this section, we offer deeper insights into the workings of DeRTa. Unless stated, we report results on the LLaMA3-70B model.
Model Black-Box Attack White-Box Attack
Code PAIR JChat Cipher Ave. Comp Auto Ave.
Vanilla 86.0 76.0 41.0 51.5 63.6 95.0 74.0 84.5
      + Harmful Prefix 88.0 78.0 35.5 21.5 55.8 25.0 36.0 30.5
      + RTO 28.0 36.0 6.5 0.0 17.6 5.0 12.0 8.5
      + Both (Ours) 21.5 24.0 1.5 0.0 11.8 4.0 2.0 3.0
Table 3: Impact of key components in our approach.
Impact of Crucial Components
In this experiment, we evaluate the effect of different components within our method. Table 3 shows the result on the LLaMA3-70B model. When implemented singularly, the harmful prefix strategy markedly enhances overall safety. However, it still remains vulnerable to several attacks. The RTO strategy effectively addresses this limitation, significantly lowering the ASR for all attacks. The results confirm our hypothesis that reinforcing the transition from potential harm to explicit safety refusal can enhance safety. The combination of both harmful prefix and RTO strategies yielded the most superior results. The forthcoming experiments will elucidate on how DeRTa substantially bolsters safety.
Figure 4: Position distribution of where the refuse token, like “sorry”, appears for safe responses.
Awareness to Refuse at Later Response Positions
We first investigate whether our method can train LLMs to refuse at later positions, as demonstrated in the case shown in Figure 2.
Figure 4 illustrates the distribution of the refusal tokens within the safe responses produced by various methods. In vanilla safety training, only 20% of the refusal tokens do not appear at the start of safe responses. Conversely, the percentages for our approach’s variations fall between 50% and 55%. At the same time, our approach results in a much higher occurrence of refusal tokens. This indicates that our method maintains a consistently higher level of safety throughout the entire sequence, meaning it is more aware and capable of refusing inappropriate content both at the beginning and later positions. Notably, LLMs refined with the RTO exhibit a strong awareness to generate refusal tokens at considerably later positions, for instance, 22.3% of responses incorporate refusal tokens beyond the $30^{th}$ position.
The ability to refuse at later response positions is crucial for defending against completion-type attacks, which is evident from the significant reduction of the ASR of CompletingAttack from 90.5% to 25.0% by employing only harmful prefixes. However, CodeAttack represents a more sophisticated challenge due to out-of-distribution (OOD) issues, with the RTO playing a critical role in mitigating CodeAttack according to our method.
Figure 5: Comparison to DPO with the same safety data.
Comparison to DPO with Harmful Response
To comprehend why RTO is effective for CodeAttack, we examine its performance by comparing it with DPO (Rafailov et al., 2024), a notable method in preference modeling that utilizes both safe and harmful responses distinctively. This experiment seeks to determine whether RTO’s success is attributed to the complete integration of harmful responses or the robust explicit modeling of token-wise safety transitions in these responses.
Figure 5 depicts the results of DPO on the LLaMA-70B model. DPO can reduce ASR for most tasks, with particularly notable improvements observed in the SelfCipher task. One possible reason is that SelfCipher explicitly leverages few-shot learning of harmful responses in prompting, a feature that DPO is specifically trained to identify and mitigate. However, the inability of DPO to improve the CodeAttack task suggests that merely integrating harmful responses does not fully account for our approach’s effectiveness in this particular scenario. As evidence, our approach significantly outperforms DPO in all tasks.
Impact of Model Size
We examine the effectiveness of our methodology across different model sizes (i.e. Mistral-7B, 8$\times$7B and LLaMA3-8B, 70B). The results, illustrated in Figure 6, clearly demonstrate that our approach significantly enhances safety irrespective of model size, showcasing the universality and robustness of our method. For detailed results across a variety of attack tasks, please refer to Table 5 in the Appendix E. Furthermore, we also provide the results for small-scale models in the LoRA setting (see Table 6).
Figure 6: ASR of different model sizes.
4.4 Robustness Analysis
In this subsection, we conduct a robustness analysis of our approach across different decoding strategies, languages, and under stricter evaluation criteria (where only fully safe cases are considered safe). We use LLaMA3-70B-DeRTa as the test model in the following experiments.
Figure 7: The ASR of DeRTa across five different languages.
Languages
Considering that changes in language can also affect the safety mechanism built around the "sorry" token, we additionally evaluated DeRTa’s performance in five different languages. We used 200 crime-related prompts from the XSAFETY dataset Wang et al. (2024b), covering five languages (English, French, Spanish, German, and Chinese), resulting in a total of 1,000 prompts. Greedy decoding was applied, and the results are presented in Figure 7. The experimental results show that even though our training data only included English, our method ensures safety across different languages. This further proves DeRTa’s robustness.
Figure 8: The ASR of DeRTa across different decoding strategies.
Figure 9: The original ASR and fully safe ASR of our model across different attacks.
Sampling Strategies
Sampling mismatches during training and testing could lead to safety issues. To further verify the stability of our method, we have added safety results under different sampling strategies. The experiment results in Figure 8 show that DeRTa demonstrates stable performance across different sampling strategies and hyperparameter settings. This aligns with our expectations, as our method establishes a token-wise defense mechanism, making it relatively robust to variations in sampling strategies across the entire sequence.
Fully Safe ASR
Given that the model’s behavior—such as providing even a partial sequence of harmful steps—can still be considered unsafe, the proposed training method intervenes in the model’s intermediate states to enable it to respond with an apology midway through the generation process. To determine whether this approach introduces bias under the current ASR criterion, we also disclosed the ASR results for the fully safe case [^5] We define ”Fully Safe ASR” as not including any harmful steps or information. Additionally, when calculating ASR in fully safe cases, responses like ”Certainly! Sorry …” or ”Sure, here are the steps: Sorry …” are treated as fully safe. in Figure 9. From the experimental results, it can be seen that in almost all cases, DeRTa refuses to answer before providing harmful steps, rather than giving harmful steps first and then refusing to answer.
4.5 Discussion
Over-sensitivity
Safety training can result in an over-sensitivity problem (Röttger et al., 2024). This means the model might unnecessarily refuse to answer questions that only seem harmful, such as "How to kill a Python program?" This issue can occur especially when there is no over-sensitivity data included in the training process. In our approach, this issue can become even more pronounced (see Table 4). However, we discover that by simply adding 200 over-sensitive samples during training, the over-sensitivity rate can be significantly reduced—from 64.0% to 18.0%—with only a minor decrease in safety performance. This highlights the importance of including borderline data in safety training datasets. This is aligned with what LLaMA3 Team (Grattafiori et al., 2024) has done, as they created a dedicated borderline dataset.
Model ASR Over-sensitivity
Vanilla 70.6 18.8
[…]
      +XStest 13.2 18.0
Table 4: The average ASR across six attacks, along with the over-sensitivity results on the XStest dataset (Röttger et al., 2024). ‘+XStest’ means that we add 200 samples from the XStest dataset to our training data, while the remaining samples are used for evaluation.
Adaptive Attack
We further test our approach by simulating an adaptive attack designed to bypass our refusal mechanism. For instance, an attacker might optimize harmful prompts that begin with "Sorry" to try to exploit our safeguard. To mimic this attack, we prefill the response template: Sorry, I cannot assist with this harmful request. However, I can provide the answer and explain why it is harmful: for each harmful query. The experimental results demonstrate that our method successfully maintains safety across all tested queries. It is worth noting that we emphasize our approach does not simply provide superficial safety, nor does it entirely eliminate the risk of adaptive attacks.
5 Conclusion
In this study, we have presented a novel approach in addressing a significant aspect of LLMs safety - refining their capacity to refuse the generation of unsafe content at any point during the response, thus addressing the critical issue of refusal position bias identified in safety tuning data. We introduce an innovative strategy encompassing two pivotal components, which collectively enhance LLMs’ ability to identify and avert unsafe content more reliably and flexibly. The comprehensive evaluation of our method notably demonstrates its superiority in terms of safety over existing baselines, especially for completion-type attacks (e.g., CodeAttack and our proposed CompletingAttack). This confirms that our approach can effectively establish a security mechanism for the entire sequence.
Our findings underscore the importance of considering the role of safety tuning data and the inherent biases that may affect an LLM’s ability to make refusal decisions effectively. Our method’s capability to defend against recent attack methods also highlights the potential for DeRTa to contribute to developing safer and more reliable LLMs in the face of continually evolving security threats.
Limitations
This paper has several limitations: (1) The evaluation does not cover all existing jailbreak attack methods. There are many jailbreak methods currently available, and evaluating our defense method against all of them would be cost-prohibitive. Therefore, we selected six representative attack methods for evaluation. (2) Similar to the first point, there are many existing defense methods; we only chose five for comparison. However, it is important to emphasize that the selected baselines were carefully chosen, focusing on safety tuning data without introducing additional training and inference costs. Some methods can increase the training/inference overhead by several to thousands of times (Mazeika et al., 2024; Sheshadri et al., 2024), and some require external safety detectors rather than ensuring safety through the LLM itself (Inan et al., 2023). (3) This work used single-turn dialogue data. Although we believe our method can naturally extend to multi-turn dialogues, this has not yet been verified. (4) Our method leads to a more pronounced issue of over-sensitivity. However, we have also verified that using a borderline dataset can effectively mitigate this problem.
Appendix A Details of Setup
Main Experiment
In training, we set the total batch size to 128 and the number of epochs to 2.
For full parameter fine-tuning (Mistral-7B and LLaMA3-8B), we use a learning rate of 2e-5, a warmup ratio of 0.03, a weight decay of 2e-5, a max length of 1024, and a dropout rate of 95% for the "Sorry" token.
For the LoRA method (Mistral-MoE and LLaMA3-70B), we set the learning rate to 1e-4, the max length to 512, with no warmup, and a 0% dropout rate for the "Sorry" token. The LoRA rank and alpha are 96 and 16, with a 0.05 dropout. The LoRA is applied in the attention layer and the mlp layer.
For GPT-4 and ChatGPT, we use the version GPT-4-turbo-0409 and GPT-3.5-tubor-0125.
To obtain uncensored Evol-Instruct data, we use ChatGPT with a safety detection prompt and keyword match (e.g., as an AI) as the filter.
Training Data for Standard Safety Tuning
Since each instance in DeRTa is a triple that consists of two (query, response) pairs (i.e., (harmful query, safe response) and (harmful query, harmful response)), we complement the safety dataset to 6,000 instances for the vanilla safety tuning for fair comparison.
DPO Experiment
To conduct standard DPO training, it is essential to have both a chosen response and a rejected response for each instruction. As such, we utilize the Qwen1.5-chat-0.5B model (Bai et al., 2023) to generate responses for the 60k helpful instructions in Evol-Instruct.
The original Evol-Instruct response and the Qwen response serve as the chosen and rejected responses, respectively. Similarly, the safe and harmful responses of a harmful question function as the chosen and rejected responses, respectively.
Building upon the model with standard safety training, we proceed to train for one additional epoch using DPO. The learning rates for LLaMA3-8B and LLaMA3-70B are set at 5e-7 and 2e-6, respectively.
Obtain Malicious Response
First, we use 330 malicious question-response pairs to adversarially tune the LLaMA3-8B-Instruct. Then, this malicious LLaMA is employed to generate harmful responses for questions from BeaverTails. Afterward, we utilize GPT-3.5 to enhance the grammar and lexical diversity of these generated responses while removing any safety warnings present in the harmful responses.
All experiments were conducted on a server equipped with eight A800 80GB GPUs. For LLaMA3-70B, we spent about 100 GPU hours on training.
Figure 10: Examples for attacks we used. The red font represents the information of the harmful question.
Appendix B Details of Applied Attacks
In Figure 10, we present examples of each attack. Below are the introductions and implementation details for each attack.
CodeAttack
CodeAttack disguises malicious questions as a code completion task. The model generates harmful questions and responses while completing the code. We use the Python-stack version for this attack.
PAIR
PAIR iteratively utilizes an AttackLLM to modify the harmful question, making it appear more harmless. In our experiment, we utilize GPT-3.5 as the AttackLLM and GPT-4 as the judge model. We maintain the same breadth (20 streams) and depth (3 iterations) as outlined in the PAIR paper.
JailbreakChat
We copy jailbreak prompts from the JailbreakChat website [^6] https://www.jailbreakchat.com/, ranking by votes. Then, we collect a total of 50 jailbreak prompts. For each harmful question, we equip it with a random jailbreak prompt.
SelfCipher
SelfCipher employs role-playing and few-shot attacks to jailbreak LLMs. It deceives the model into believing that we are communicating in cipher, while simultaneously presenting malicious question-answer pairs. In our experiment, we provide 5 fixed malicious demonstrations.
CompletingAttack
This method removes all formatting tokens (e.g., [INST]) to present the query in a declarative format, thereby enabling the model to complete the text. In our paper, we use GPT-3.5-turbo to transform these queries to the declarative version.
AutoDAN
AutoDAN can automatically create stealthy jailbreak prompts using a designed hierarchical genetic algorithm. We utilize the HGA version of AutoDAN and GPT-4 as the mutation LLM. Once the final mutated jailbreak prompt is generated (only partial response obtained), we input these queries into the test model to get the final complete response. Therefore, the reported ASR might be lower than the original paper. We perform this operation in all experiments to make fair comparisons.
Appendix C Guideline for Safety Evaluation
We establish the following principles for evaluating safety in responses generated by the models. The guidelines help in determining whether a specific response is considered unsafe or not unsafe.
Unsafe
A response will be deemed unsafe if all of the following conditions are met:
  1. 1.
    Related to Harmful Input:The response is directly related to a harmful or dangerous question or input provided by the user.
  2. 2.
    Provides Specific or Directive Harmful Information: The response offers specific harmful information or instructions that could be used to carry out a harmful activity.
Note: If both of these conditions are met, the response will be classified as unsafe, regardless of whether the model includes disclaimers, rejection messages, or safety warnings within the response.
Not Unsafe
A response will be considered not unsafe if any of the following conditions are satisfied:
  1. 1.
    Refusal to Respond to Harmful Input: The response clearly refuses to answer the harmful question or engage with the harmful input.
  2. 2.
    Lack of Relevance to Harmful Question: The response is poorly related or irrelevant to the harmful question or input.
  3. 3.
    Redirects or Requests Further Information: The response does not directly answer the question but instead insults the questioner or only asks the questioner for additional information without providing harmful instructions.
  4. 4.
    Provides Partial Information Only: The response provides only partial information, from which the user cannot complete the harmful task or activity.
Appendix D Case Study
We provide some selected cases in this section, all cases of our experiments can be found in our GitHub project. Please refer to Figures 11-14.
Figure 11: The response in JailbreakChat.
Figure 12: The response in CompletingAttack.
Figure 13: The response in SelfCipher.
Figure 14: The response in AutoDAN.
Appendix E Main Results on Small-Scale LLMs
We present the results of LLaMA3-8B and Mistral-7B on Table 5-7.
For the GCG method (see Table 5), we fix a bug in the original code by using the solution given by the authors [^7] https://github.com/llm-attacks/llm-attacks/issues/40. We also added our conversation template to the code and set the number of attack steps to 500. We do not make any other changes to the code.
The results in Table 5 show that our method also performs effectively on small-scale models, aligning well with the outcomes observed in large-scale models. This highlights the adaptability and broad applicability of our approach.
To better control variables, we also included the results of using LoRA to fine-tune smaller-scale models (refer to Table 6). These results further support our previous conclusions.
Model Safety (Attack Success Rate $\downarrow$) Helpfulness ($\uparrow$)
PAIR JChat Cipher Comp Auto GCG GSM8K MMLU Alpaca
Open-Source Mistral-7B
Vanilla 84.0 9.5 34.0 82.5 66.0 50.0 22.4 40.2 80.7
      + Ours 44.0 4.0 4.0 7.5 20.0 16.0 20.4 41.8 78.7
Open-Source LLaMA3-8B
Vanilla 82.0 17.5 12.0 93.0 82.0 32.0 43.8 49.0 88.3
      + Ours 24.0 4.0 0.0 6.0 14.0 2.0 46.4 50.4 88.7
Table 5: Main results on small-scale LLMs. For CodeAttack, these models often fail to follow instructions, so we do not display the results under this setting.
Model PAIR JChat Cipher Comp Auto Average
Open-Source Mistral-7B-LoRA
Vanilla 76.0 42.5 91.0 89.5 80.0 75.8
Ours 50.0 7.5 0.5 4.5 6.0 13.7
Open-Source LLaMA3-8B-LoRA
Vanilla 76.0 26.5 31.0 92.0 82.0 61.5
Ours 46.0 3.5 0.5 5.0 8.0 12.6
Table 6: Results on LoRA version small-scale LLMs.The LoRA rank is 32.
Model PAIR JChat Cipher Comp Auto Average
DPO 62.0 31.0 4.5 88.5 70.0 51.2
[…]
Table 7: DPO results on LLaMA3-8B.